OpenStudy (anonymous):

whats an equation for the line having (1,-6)and(-4,4)?

OpenStudy (anonymous):

m=(4-(-6))/(-4-1)=10/(-5)=-2 this is the slope of your line. use formula for the line: y-(-6)=-2*(x-1) y+6=-2x+2 y=-2x-4